OpenStudy (anonymous):

Solve for x: -5(x - 3) = 30

OpenStudy (anonymous):

-3 medal please

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

-3-3=-6. -5*-6=30. good job

OpenStudy (anonymous):

What is the equation of the line that passes through the points (-3, 2) and (-5, 8)? 3x - y = -11 3x + y = - 7 3x + y = 5 x + 3y = 3